OpenStudy (liizzyliizz):

dy/dx of y^3*x+y^2*x^2 = 6 I keep getting this wrong by hand, I have a feeling I know what I am doing wrong but can someone go through it step by step?

OpenStudy (anonymous):

you need to use implicit differentiation yes I will help you step by step

OpenStudy (anonymous):

show me what you have so far and I will help you where you go wrong sound good?

OpenStudy (liizzyliizz):

sure, I know for a fact I am forgetting the x for the y's when you have dy/dx but I honestly forgot how to do that lol. so I'll do ti by hand now :p

OpenStudy (anonymous):

Okay =)

OpenStudy (anonymous):

just a tip it is a lot easier if instead of writing dy/dx you just write y' and x' for y prime and x prime it just makes it easier on yourself to read you know what I mean?

OpenStudy (anonymous):

satellite I got this

OpenStudy (anonymous):

ok and yes, your tip is a good one i was just going to make a small suggestion

OpenStudy (liizzyliizz):

yeah I used to do that lol, but since I have my AP exam tomorrow I want to get into the habit of using dy/dx so I won't confuse myself or a grader if it happens to be found in the free response lol.

OpenStudy (anonymous):

alright fair enough =)

OpenStudy (anonymous):

sorry satellite didn't mean to sound mean I just like the problems where I get to help the person through it haha.

OpenStudy (anonymous):

just that if it was not clear how to work the with the product rule and chain rule in these, imagine that \(y=f(x)\) or maybe even \(y=\sin(x)\) and then see how you would find the derivative of \[\sin^3(x)*x+\sin^2(x)*x^2 = 6 \] sometimes this clarifies the method

OpenStudy (anonymous):

liizzyliizz do you know any implicit differentiation or do you want me to work you towards the problem with easier examples?

OpenStudy (liizzyliizz):